A ray of light passing grom water to air is deviated 10° from its original direction.what is the a) angle o incidence b) angle of reflection c) angle of refraction?


1
Expert's answer
2022-04-25T13:33:14-0400

"\frac{\sin \alpha}{\sin(\alpha+10)}=\frac 1n=0.75,\implies"

"\alpha=26.5°,"

"\beta=\alpha=26.5°,"

"\gamma=36.5°."